Please do 2,3 and 4
Experiment 3 Pr 1. A student weigh be 98.595 oth mark, the stoppered flask is weighed again and the mass is found to be 204.274 & ment s an empty 100 mL. volumetric flask and stopper and finds the mass to g After adding 4.163 g of sugar to the flask, dissolving it, and filling the flask Calculate the density of this solution dm oS 629 2. Explain how to properly prepare a solution in a 100 mL volumetric flask. accidentally add too much water, and the level rises above the mark. You decide to remove the excess water with a pipet, and then you measure the mass of the stoppered flask plus solution. Explain why the density calculated using this data will not be accurate. 3. Suppose that in adding water to a volumetric flask containing dissolved sugar, you
media%2Fba4%2Fba47e85b-e141-40c2-912e-ce
Answer #1

(2) Weigh the solid in a glass weighing bottle.

Transfer the solid to a volumetric flask.

Add water/solvent to volumetric flask to halfway mark.

Rinse the weighing bottle with the water/solvent and transfer to the flask.

Stir/shake the solution to dissolve the solid.

Make up the volume with water/ solvent up to the mark of 100 mL.

(3) Once you added an excess of water, a small portion of sugar might already be dissolved in the excess water, so If you remove a small/excess portion the little-dissolved sugar will also get lost hence the measured density will be less.

(4) Volume = 100 mL

Mass of solution = 103.87 g

Density = 103.87 g / 100 mL = 1.0387 g/mL

The fitted linear equation, y = mx + c

slope = m = 0.0038

intercept = c = 0.9959

Density = y = 1.0387 g/mL

Mass of sugar, x = (y - c) / m = (1.0387 - 0.9959) / 0.0038 = 11.2632 g